• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Ransome

    Log entry time 07:58:32 on December 9,2008

    Entry number 252677

    keyword=Shift Summary

    Begin shift with run 4308 running
    00:35 - stop run 900 K
    Brad Sawatzky will do some tests with Big Bite
    1:30 - Brad done, start 4332
    2:40 - 4332 ended - 1 million
    2:41 - begin 4333 -
    3:45 - end 4333 1 million
    3:46 - start 4334
    4:50 - end 4334 1 million
    4:51 - start 4335
    5:35 MCC calls - they are having problem with laser in injector
    5:50 stop run 4335 - 620 K
    MCC calls - problem with injector - expert called
    6:45 - MCC will send tuned beam
    6:50 - MCC sends 2 uA - we have empty target
    checked rates with 2 uA and 3x4 and 5x6.6 - look the same
    7:10- 4337 started - 8 uA - MCC reports still some problems, beam
    may be unstable.
    current remains at about 7 uA through end of shift. No word from MCC